OpenStudy (anonymous):

define the order of operations. @mehek14

OpenStudy (anonymous):

The rules that say which calculation comes first in an expression They are: • do everything inside parentheses first: () • then do exponents, like x2 • then do multiplies and divides from left to right • then do the adds and subtracts from left to right Example: 5 × (3 + 4) − 2 × 8 = 5 × 7 − 2 × 8 = 35 − 16 = 19

OpenStudy (anonymous):

Like Pemdas? Parenthesis Exponents Multiplication Division Addition Subraction

Nnesha (nnesha):

order of operation \(\huge\color{green}{{\rm PEMDAS}}\) Parentheses Exponents MD multiply or divide AS add or subtract first you have to solve parentheses then exponent multiply or divide and last subtract or add

OpenStudy (anonymous):

thx :)